# UniRepLKNet: A Universal Perception Large-Kernel ConvNet for Audio, Video, Point Cloud, Time-Series and Image Recognition

### 27 Nov 2023 v1

@@ 6 Conclusion
We proposed four architectural guidelines for large-kernel ConvNet and an architecture named UniRepLKNet, which not only shows a leading performance in image recognition but also achieves remarkable results even on modalities such as audio and time-series data, outperforming multiple specialist models on those modalities. Such results not only signify a comeback for ConvNet in its original domain but also showcase large-kernel ConvNet’s potential to conquer new territories, highlighting their adaptability and broad utility across different modalities and tasks.
@@
We present the detailed training configurations for image classification, object detection, and semantic segmentation. We will publicly release a reproducible training script and trained weights for every model.
@@
COCO object detection. For fair comparisons, we follow common practices [37, 40] to initialize the backbone with pretrained weights and train the models using a 3$\times$ (36 epochs) schedule by default. The shorter side is resized to 480--800 pixels, while the longer side does not exceed 1,333 pixels. All the models are trained with a batch size of 16 and AdamW [41] optimizer with an initial learning rate of $1\times 10^{-4}$. We will publicly release the training configuration file used in the MMDetection framework.
ADE20K semantic segmentation. We evaluate UniRepLKNet models on the ADE20K dataset [79], and initialize them with the pre-trained classification weights. The learning rate is initialized with $1\times 10^{-4}$ and decayed with the polynomial decay schedule with a power of 1.0. Following previous methods [37, 40], the crop size is set to 512 for the ImageNet-1K-pretrained models, and 640 for ImageNet-22K-pretrained models. All segmentation models are trained with a batch size of 16 for 160k iterations. We will publicly release the training configuration file used in the MMSegmentation framework.

### 18 Mar 2024 v2

@@ 6 Conclusion
UniRepLKNet shows a leading performance in image recognition and achieves remarkable results even on modalities such as audio and time-series data, outperforming multiple specialist models on those modalities. Such results signify a comeback for ConvNet in its original domain and showcase large-kernel ConvNet’s potential to conquer new territories. The limitations are noticeable, e.g., the dilated branches require more training resources, which may be upgraded with simpler [5] or gradient [18] re-parameterization; the applications to large vision-language models [38, 45, 79], cross-attention-based scenarios [96, 7], and generation tasks [58, 95] remain under-explored.
Acknowledgements. This work is partially supported by the National Natural Science Foundation of China (Grant No. 8326014).
@@
We present the detailed training configurations for image classification, object detection, and semantic segmentation. We have publicly released a reproducible training script and trained weights for every model on GitHub.
@@
COCO object detection. For fair comparisons, we follow common practices [49, 52] to initialize the backbone with pretrained weights and train the models using a 3$\times$ (36 epochs) schedule by default. The shorter side is resized to 480--800 pixels, while the longer side does not exceed 1,333 pixels. All the models are trained with a batch size of 16 and AdamW [53] optimizer with an initial learning rate of $1\times 10^{-4}$. We have publicly released the training configuration files used in the MMDetection framework and trained weights.
ADE20K semantic segmentation. We evaluate UniRepLKNet models on the ADE20K dataset [97], and initialize them with the pre-trained classification weights. The learning rate is initialized with $1\times 10^{-4}$ and decayed with the polynomial decay schedule with a power of 1.0. Following previous methods [49, 52], the crop size is set to 512 for the ImageNet-1K-pretrained models, and 640 for ImageNet-22K-pretrained models. All segmentation models are trained with a batch size of 16 for 160k iterations. We have publicly released the training configuration files used in the MMSegmentation framework and trained weights.
Appendix C: Shape Bias
Figure 6: Shape bias of ImageNet-22K-pretrained UniRepLKNet-L and RepLKNet-31L.
Figure 7: Shape bias of ImageNet-1K and ImageNet-22K-pretrained RepLKNet-31B and Swin-B. This figure is directly taken from the supplementary material of RepLKNet without any modifications
A higher shape bias means the model makes predictions based more on the shape of objects rather than the textures, i.e., the model behaves more similarly to humans. Therefore, a model with a higher shape bias may transfer better to downstream tasks. UniRepLKNet demonstrates significantly higher shape bias than existing ConvNets and ViTs. Concretely, we test the shape bias of ImageNet-22K-pretrained UniRepLKNet-L and RepLKNet-L with the modelvshuman toolbox  [^5] https://github.com/bethgelab/model-vs-human. Fig. 6 shows a significantly higher shape bias of UniRepLKNet - UniRepLKNet makes 20% more decisions based on the overall shapes of objects. This improvement is particularly remarkable since RepLKNet is already known to have a high shape bias (Fig. 7 is directly taken from the supplementary material of the RepLKNet paper without any modifications).
6.1 Appendix D: Training Memory Footprint
The extra parallel dilated branches in Dilated Reparam Block consume more training resources, which is acceptable considering the performance improvements. We present the peak GPU memory footprint and training speed in Table 16. With a bigger model and bigger data, we may trade the performance for higher training speed and lower memory consumption by replacing the Dilated Reparam Block with a single large-kernel conv layer followed by Batch Normalization layer. We test the peak memory footprint and actual training throughput while training UniRepLKNet-S with 224$\times$224 inputs and a batch size of 4096 on a node with eight A100 GPUs. Note that such results are significantly influenced by the hardware environment and specific implementation; thus, they should be considered as references only.
Table 16: Training costs.
Peak memory Training throughput
Dilated Reparam Block 24.6GB 6642 images/s
Single large-kernel conv layer 20.8GB 9675 images/s